# Ambiguous Resonances in Multipulse Quantum Sensing with Nitrogen Vacancy Centers

### v2

@@ VI Finite Pulse Duration
In most applications of multipulse sequences, perfect $\delta$ pulses with zero temporal length are assumed. Even though the pulse length might still be much smaller than the pulse separation $t_{\pi}\ll\tau$, it has been shown that the free evolution of the system during finite pulse durations under the influence of $H_{0}$ leads to additional resonances at even multiples of the resonant frequencies and their odd subharmonics [16], that is $f=2k/(2l+1)f_{0}$ for $k$ and $l$ natural. Notably, lattice 13C nuclear spins can give rise to ambiguous resonances of 1H, other than the ones already described in the previous section. This occurs because the 4th harmonic of a weakly coupled 13C coincides with the fundamental resonance of 1H, given that $42.824\gamma^{c}=$ 42.82 MHz/T and $\gamma^{h}=42.57$ MHz/T. These ambiguous resonances can still be greatly suppressed by adding a random phase $\Phi_{m}$ for each XY8m block, known as the RXY8-$M$ sequence [2, 45]. A further improvement of this suppression was theoretically proposed by correlating the random phases such that they cancel each other every $g$ repetitions of the XY8 block, $\sum_{m}^{g}\exp(-i\Phi_{m})=0$ [3].
@@
Figure 6: XY8-12 and RXY8-12 sequences with resonance from an external RF signal with $f_{0}=5.5$ MHz for $m_{s}=0\leftrightarrow-1$ transition. Multiple fringes are observed around the fundamental resonance $\tau_{0}=1/2f_{0}$, in addition, spurious harmonics at $\tau_{0}/\tau=$ 3/4, 5/4, 3/2 and 7/4 are suppressed by the phase randomization. Random phase correlations with $g=2$ and 3 are also demonstrated, without much differences from uncorrelated RXY8. The best correlations $r=$(0.839, 0.877, 0.818, 0.888) occur at $\gamma_{e}B_{2}^{z}=0.28$ MHz and $\gamma_{e}B_{2}^{x}=\gamma_{e}B_{2}^{y}=0$ from Eq. 4.
@@
To experimentally reproduce these ambiguous resonances, an external RF signal with frequency $f_{0}=5.5$ MHz was applied to a single NV at bias field $B_{0}=38$ mT and $\pi$-pulse length $t_{\pi}=17.16$ ns. Four different multipulse sequences were applied as shown in Fig. 6. In all four we observe multiple fringes around the fundamental resonance $\tau_{0}=1/(2f_{0})$, resulting from the convolution of the spectral density of the RF field and the filter function of the multipulse sequence [31]. Apart from that, XY8-12 has four of these ambiguous resonances at $f/f_{0}=$ 2/3, 4/3, 4/5 and 4/7, which are already greatly suppressed with RXY8-12. We also demonstrate the experimental realization of RXY8-12 with correlated phases at $g=2$ and $g=3$ repetitions, with no noticeable differences from uncorrelated RXY8 at these experimental conditions. Another advantage of introducing the random phase $\Phi_{m}$ lies in the robustness against pulse errors [3], such as in the pulse length and detuning of the MW resonant frequency. These lead to a degradation of the resonance contrasts and are discussed in more details in SM:V.
@@
VII Discussion and Simulations Dataset
@@
Noise filtering by DD sequences has proven to be an effective technique in quantum sensing, capable of detecting very weak oscillating fields of even single nuclear spins. This great sensitivity comes at the cost of an often complex signal analysis influenced by external and inherent interactions which lead to ambiguous resonances, such as the ones studied in this paper. It is therefore imperative that these effects are well understood and characterized for an accurate identification of nuclear spin resonances. Among candidate platforms, the NV center in diamonds outstands in its ease of implementation and versatility with different biological and chemical systems. Nevertheless, widely adopted approximations and oversimplifications of the system dynamics completely overlook these ambiguous resonances. Examples of such approximations are the RWA, assuming perfect $\delta$-pulses, adopting a reduced electronic spin manifold where $S=1/2$ or ignoring the nitrogen nuclear spin of the center.
In this paper, we proposed a numerical simulations model for the time-dependent dynamics of the NV under a DD sequence and experimentally characterized three effects that lead to ambiguous resonances. Where in all three cases, the numerical simulations model showed robust correlations with the experimental data. Regarding 15NV, a small misalignment of the bias magnetic field gives rise to non-commuting elements with the rotation operator of the RWA. Which in turn, can cause the precession of 15N nuclear spins to be sensed via the hyperfine interaction. These resonances are present for pulse separations $\tau>250$ ns and differ depending on the $m_{s}$ state. The small sensitivity of the resonant frequencies with the magnetic field $B_{0}$ evidences the dominance of the hyperfine interaction with the electronic spin, over the Zeeman interaction with $B_{0}$. We have also shown that in the presence of an external misaligned field around 40 mT, the 15N resonances closely resemble those from 1H nuclear spin for different XY8-$M$ orders. This characterization of 15NV resonances provides a basis for future studies aimed at using the 15N nuclear spin as a quantum memory [6, 40].
Another more studied source of ambiguous resonances are the diamond lattice 13C nuclear spins, which can also couple to the NV electronic spin with resonant frequencies depending on the hyperfine interaction matrix $A^{c}$. Here, we have presented a method for unequivocally calculating $A^{c}$, based on the Hahn echo decay modulation frequencies as a function of the bias field $B_{0}$. Followed by a comparison with numerically calculated eigenvalues from the total Hamiltonian. The method is further validated by the XY8-$M$ measurements and simulations at different fields and orders. This also demonstrates that the numerical simulation model can be applied to higher dimensional Hilbert spaces, in this case including the sensed spin. Although applied for 13C nuclear spins, the method is general to any coupled spins in the strong coupling regime [9], as long as they can be described in terms of a Hamiltonian as in Eq. 3.
Specific parameters of the MW control pulses which are not accounted for in perturbation theory also affect the resonances of DD sequences. In particular, the free evolution of the probing spin during finite pulse durations leads to resonances at the even multiples of the fundamental resonant frequencies [16]. These resonances, however, can be suppressed with a phase randomization in each XY8 block [2, 45]. In this paper, we experimentally demonstrated the RXY8 with random phase correlation, which has been proposed to further improve spurious harmonics suppression and robustness against pulse errors [3]. The capability of the simulations model to account for the specific time-dependency of the control field $H_{1}(t)$, Eq. 2, further underlines its versatility.
@@
Unlike the ambiguous resonances originating from the finite pulse durations, a straightforward method for suppressing the resonances from 15N and 13C in DD sequences remains an open problem. Currently, this would involve more elaborate experimental conditions, as precise vector magnets, isotopically purified diamonds or more complicate pulse sequences. Therefore, some of the ambiguous resonances presented here are unavoidable in a large variety of experiments. A platform for analyzing and identifying spectral features is thus essential for the continued development of the field, as these ambiguous resonances are largely dependent on specific experimental parameters, resulting in numerous possible resonant responses in multipulse DD sequences.
@@
To further extend the toolbox of available methods, we provide a dataset comprised of around $10^{4}$ simulations of XY8-$M$ for 15N and 13C ambiguous resonances. Both $m_{s}$ transitions were considered with several different combinations of $B_{0}$ and order $M$. It is important to note however, that a fixed Rabi frequency of $\gamma_{e}B_{1}=40$ MHz was considered in all simulations, which may lead to slight differences at high order XY8. For 15N resonances, the angle $\theta_{0}$ was also changed, whereas for 13C, the hyperfine matrices of all provided families [41] were considered. The simulations can be visualized and compared with the user’s experimental data for spectral disambiguation through a user-friendly graphical interface, published under DOI:10.6084/m9.figshare.26245895 [47], where a complete guide for installation and usage are also provided. Even though the dataset only contains XY8-$M$ simulations for NVs so far, it can be used for other DD sequences which have similar resonance spectra, such as the CPMG [13]. Furthermore, the model and results here presented have the potential to be extended to other color centers or quantum systems in general, as well as to more elaborate DD sensing protocols.
@@ Acknowledgements.
Supplementary Materials: Ambiguous Resonances in Multipulse Quantum Sensing with Nitrogen Vacancy Centers
S1 Experimental Setup
@@
The experiments were performed on a custom-made confocal microscope. A diode laser with a wavelength of 518 nm (iBEAM-SMART-515-S, TOPTICA Photonics AG) was used for both continuous and pulsed optical excitation of an NV center. Emitted fluorescence was collected either with an air objective (MPlanApo N 50$\times$/0.95, Olympus) with NA = 0.95 or an oil objective (HCX PL APO 100$\times$/1.40, Leica; type F immersion liquid, Leica) with NA = 1.40 and detected by an avalanche photodiode (SPCM-AQRH-14, Excelitas). The laser light was blocked from entering the detector by a 650 nm longpass filter (FELH0650, Thorlabs). Continuous microwave field for ODMR spectra was generated by a microwave source (TGR6000, TTi) and the radiofrequency signal in Sec. IV was generated by a function generator (TGF4242 LXI, TTi). Microwave pulses for dynamical decoupling multipulse sequences were generated by an arbitrary waveform generator (AWG7122C, Tektronix). In both cases generated microwave fields were amplified (ZHL-16W-43+, MiniCircuit) and then transmitted in the NV centers’ proximity by a copper (99.99%) wire with a diameter of 20 $\mu$m. Magnetic field was applied with a permanent magnet mounted on a holder with 4 degrees of freedom and with the maximum achievable magnetic field projection on NV center axis of 0.1 T. For time-resolved photon counting we used a time-correlated single photon counter (MCS6, FAST ComTec).
S2 Diamond Samples
@@
In Sec. IV of the main text, the effects of 15N coupling on Hahn echo were measured in a IIa diamond with a 12C overgrown layer of 90 nm, which was later implanted with 15N at energy 2 keV and annealed at 1000C in ultra-high vacuum. For the XY8-$M$ measurements, another shallow implanted sample was used with implantation energy 2 keV and dose $10^{9}$ 15N+/cm2, annealed at 10-7 mbar and 1000C. In Sec. V, the 13C ambiguous resonance were studied in an electronic grade diamond grown by chemical vapor deposition, with natural abundance of 13C and 14N, in order to avoid 15N effects. Finally, the ambiguous resonances from finite pulses in Sec. VI were measured in a thin electronic grade diamond plate with deeper 15N implantation (20-30 keV) and doses around $10^{9}$ 15N+/cm2, also annealed in ultra high vacuum at 1000C.
S3 Pearson Correlation Coefficient
As discussed in the main text, experimental fluorescence intensity and simulated transition probability to the bright state are related through a linear equation. Therefore, an appropriate metric to determine the correlation between both is the Pearson coefficient $r$ [1], which is the square root of the coefficient of determination $R^{2}$ for linear relations. It is defined as the ratio between the covariance of the two data sets and the product of their standard deviations, where one data set represents the experimental data and the other the simulation. Within this definition, $r$ is normalized from -1 to 1, with $r=0$ representing no linear dependency between experimental and simulated data, while $r=1$ indicates a perfect linear correlation.
To illustrate this, Fig. S1 shows simulations for XY8-2 of the 15N ambiguous resonances at $B_{0}=17$ mT for different pulse length deviations in regards to the actual $\pi$-pulse duration $T_{p}/t_{\pi}$ (Sec. IV of main text). Compared to the actual experimental data, assuming $T_{p}/t_{\pi}=1$ results in a low correlation of $r=0.555$ and thus an inadequate simulation of the experimental data, whereas taking $T_{p}/t_{\pi}=0.826$ provides a better correlation of $r=0.858$. In this case, the pulse length deviation affects the relative intensity of each resonance peak keeping their $\tau$ fixed, while also inducing a high frequency modulation of the signal.
@@
Figure S1: The XY8-2 of the ambiguous 15N resonance (Sec. IV of main text) as a function of the pulse duration $T_{p}/t_{\pi}$ exemplifies the use of the Pearson coefficient, with r = (0.419, 0.858, 0.555, 0.685) for $T_{p}/t_{\pi}=$ (0.620, 0.826, 1.000, 1.215) respectively. At $T_{p}/t_{\pi}=1$ the simulated and experimental data have a poor correlation, but considering a small deviation, the experimental data can be better modeled by the simulation.
S4 13C Hahn Echo Modulations
@@
The Hahn echo signal at $B_{0}=31$ mT and $B_{0}=18$ mT for both $m_{s}$ transitions are shown in Fig. S2. As can be observed, the simulations and experimental data both present the same modulation frequencies, however the amplitude of the fast modulation is smaller in the experiments. This could be related to some experimental error or some unconsidered factor in the Hamiltonian model. Nonetheless, this discrepancy does not affect the study of the ambiguous resonances and as seen in the main text, the calculated hyperfine matrix and the numerical model still reliably described them.
@@
Based on those experimental values of the modulation frequencies, over $10^{7}$ possible combinations of hyperfine coupling matrices between the NV and the 13C nuclear spin were tested in the total Hamiltonian $H_{0}+H_{2}$ from Eqs. 1 and 3 of the main text. The eigenenergies of the Hamiltonians were calculated numerically, with the effective Larmor frequencies $\tilde{\omega}^{c}_{m_{s}}$ corresponding to the energy differences between the carbon nuclear states in each electronic spin states $m_{s}$. The hyperfine matrix with the least square difference in units of MHz was
@@
Figure S2: Hahn echo modulations from a 14NV coupled to a 13C nuclear spin, resulting in correlations between experimental and simulated data of $r=$ (0.925, 0.792) at $B_{0}=31$ mT and $r=$ (0.806, 0.659) at $B_{0}=18$ mT. Although both agree in regards to the modulation frequencies, the amplitude of the faster modulation is smaller experimentally, which could be due to some experimental error or unconsidered factor in the Hamiltonian.
S5 Pulse Errors
Phase randomization can greatly increase robustness against pulse errors [2, 3]. In non-randomized XY8, small detunings in the MW frequency $\delta_{0}$ and deviations in the pulse length $T_{p}/t_{\pi}$ lead to a degradation of the resonance contrast and additional peaks over the $\tau$ range, which could represent another ambiguous resonance source for similar frequency RF fields or spins. To show this effect, XY8-2 and RXY8-2 with uncorrelated phases were measured and simulated with an applied RF field of $f_{0}=5$ MHz, taking $\delta_{0}=20$ MHz and $T_{p}/t_{\pi}=0.7$ separately (Fig. S3). Experimental parameters are $t_{\pi}=15.24$ ns and $B_{0}=34$ mT with $\theta_{0}=4.1^{\circ}$. Again, we observe a suppression of the spurious peaks in RXY8-2 and a further validation of the simulation model. In this case at low order $M=2$, the RF signal can be approximated by an effective hyperfine interaction with $A^{s}_{xz}=0.92$ MHz and $\gamma^{s}B_{0}=5$ MHz, as in Eq. 4 of main text. Whereas for the higher order XY8-12 from Sec. VI of the main text, the quantum mechanical approximation of the classical RF field already deviates from the experimental data.
Figure S3: XY8-2 and RXY8-2 with external RF signal of $f_{0}=5$ MHz with MW pulse deviation of $\delta_{0}=20$ MHz or length error of $T_{p}/t_{\pi}=0.7$. The RF field is approximated by an effective hyperfine coupling with $A^{s}_{xz}=0.92$ MHz, resulting in $r=$(0.763, 0.907, 0.626) for XY8-2 and $r=$(0.639, 0.915, 0.423) in RXY8-2. The pulse errors lead to a degradation of the contrast and additional resonances, which are suppressed by the random phases.

### v3

@@ IV.3 Finite Pulse Duration
In most applications of multipulse sequences, perfect $\delta$-pulses with zero temporal length are assumed. Even though the pulse length might still be much smaller than the pulse separation $t_{\pi}\ll\tau$, it has been shown that the free evolution of the system during finite pulse durations under the influence of $H_{0}$ leads to additional resonances at even multiples of the resonant frequencies and their odd subharmonics [13], that is $\tau=(2k+1)/(2l)\tau_{0}$ for $k$ and $l$ natural. Notably, lattice 13C nuclear spins can give rise to ambiguous resonances of 1H, other than the ones already described in the previous section. This occurs because the 4th harmonic of a weakly coupled 13C coincides with the fundamental resonance of 1H, given that $42.824\gamma^{c}=$ 42.82 MHz/T and $\gamma^{h}=42.57$ MHz/T. These ambiguous resonances can still be greatly suppressed by adding a random phase $\Phi_{m}$ for each XY8m block, known as the RXY8-$M$ sequence [49, 50]. A further improvement of this suppression was theoretically proposed by correlating the random phases such that they cancel each other every $g$ repetitions of the XY8 block, $\sum_{m}^{g}\exp(-i\Phi_{m})=0$ [51].
@@
Figure 5: XY8-12 and RXY8-12 sequences with resonance from an external RF signal with $f_{0}=5.5$ MHz for $m_{s}=0\leftrightarrow-1$ transition. Multiple fringes are observed around the fundamental resonance $\tau_{0}=1/2f_{0}$, in addition, spurious harmonics due to the finite pulse length at $\tau/\tau_{0}=$ 3/4, 5/4, 3/2 and 7/4 are suppressed by the phase randomization [49, 50]. A proposed improvement of the RXY8 sequence by random phase correlations with $g=2$ and 3 [51] is also demonstrated, without significant differences from uncorrelated RXY8. Thus, indicating that phase correlation can be introduced in the RXY8 sequences, without affecting non-ambiguous spectral features. The best correlations $r=$(0.839, 0.877, 0.818, 0.888) occur at $\gamma_{e}B_{2}^{z}=0.28$ MHz and $\gamma_{e}B_{2}^{x}=\gamma_{e}B_{2}^{y}=0$ from Eq. 4.
@@
To experimentally reproduce these ambiguous resonances, an external RF signal with frequency $f_{0}=5.5$ MHz was applied to a single NV at bias field $B_{0}=38$ mT, where the MW $\pi$-pulse length was $t_{\pi}=17.16$ ns. Four different multipulse sequences were applied as shown in Fig. 5. In all four we observe multiple fringes around the fundamental resonance $\tau_{0}=1/(2f_{0})$, resulting from the convolution of the spectral density of the RF field and the filter function of the multipulse sequence [30]. Apart from that, XY8-12 has four of these ambiguous resonances at $\tau/\tau_{0}=$ 3/4, 5/4, 3/2 and 7/4, which are almost entirely suppressed with RXY8-12. Furthermore, we demonstrate the experimental realization of RXY8-12 with correlated phases at $g=2$ and $g=3$ repetitions. At these experimental conditions, the two sequences show no noticeable differences from uncorrelated RXY8. Meaning that they can be used instead of the RXY8, as they have additional robustness against pulse errors [51], such as in the pulse length and detuning of the MW resonant frequency. These lead to a degradation of the resonance contrasts and are discussed in more details in Appendix E.
@@
V Discussion and Simulations Dataset
@@
Noise filtering by DD sequences has proven to be an effective technique in quantum sensing, capable of detecting very weak oscillating fields of even single nuclear spins. This great sensitivity comes at the cost of a complex signal analysis influenced by external and inherent interactions which lead to ambiguous resonances, such as the ones studied here. It is therefore imperative that these effects are well understood and characterized for an accurate identification of nuclear spin resonances. Among candidate platforms, the NV center in diamonds outstands in its ease of implementation and versatility with different biological and chemical systems. Nevertheless, widely adopted approximations and oversimplifications of the system dynamics completely overlook these ambiguous resonances. Examples of such approximations are the RWA, assuming perfect $\delta$-pulses, adopting a reduced electronic spin manifold where $S=1/2$ or ignoring the nitrogen nuclear spin of the NV.
In this work, we presented a numerical simulations model for the time-dependent dynamics of the NV under a DD sequence and experimentally characterized three effects that lead to ambiguous resonances. First, in 15NV centers, a small misalignment of $\vec{B}_{0}$ causes the precession of 15N nuclear spins to be sensed via the hyperfine interaction. As demonstrated here, these resonances can closely resemble a 1H nuclear spin. Another relevant source of ambiguous resonances are the diamond lattice 13C nuclear spins, which also couple to the NV electronic spin through the hyperfine interaction matrix $A^{c}$. Here, we have presented a method for unequivocally calculating $A^{c}$, based on the ESEEM frequencies as a function of the bias field $B_{0}$. This method can be generalized to any coupled spins in the strong coupling regime [9], which can be expressed in terms of a Hamiltonian as in Eq. 3. Lastly, specific parameters of the MW control pulses which are not accounted for in perturbation theory also affect the resonances of DD sequences. In particular, the free evolution of the probing spin during finite pulse durations leads to resonances at the even multiples of the fundamental resonant frequencies [13]. These resonances can be suppressed with phase randomization in each XY8 block [49, 50], where we experimentally demonstrated a proposed improvement with random phase correlation [51].
@@
Unlike the ambiguous resonances originating from the finite pulse durations, a straightforward method for suppressing the resonances from 15N and 13C remains an problem. Currently, this would involve more elaborate experimental conditions, as precise vector magnets, isotopically purified diamonds or more complicate pulse sequences. Another important technique that can be potentially employed for improving pulse fidelity and compensate for finite-length hard pulses is pulse shaping [52], which could be easily incorporated in our numerical simulations model. In conclusion, some of the ambiguous resonances presented here are simply unavoidable in a large variety of experiments. A platform for analyzing and identifying spectral features is thus essential for the continued development of the field, as these ambiguous resonances are largely dependent on specific experimental parameters, resulting in numerous possible resonant responses in multipulse DD sequences.
@@
To further extend the toolbox of available methods, we provide a dataset comprised of around $10^{4}$ simulations of XY8-$M$ for 15N and 13C ambiguous resonances. Both $m_{s}$ transitions were considered with several different combinations of $B_{0}$ and order $M$. It is important to note however, that a fixed Rabi frequency of $\gamma_{e}B_{1}=40$ MHz was considered in all simulations of the dataset, which may lead to slight differences at high order XY8. This typically high Rabi frequency was chosen in order to mitigate the ambiguous resonances from finite pulse durations and thus not overlap it with the two other sources of ambiguous resonances. For 15N resonances, the angle $\theta_{0}$ was changed, whereas for 13C, the hyperfine matrices of all provided families [46] were considered. The simulations can be visualized and compared with the user’s experimental data for spectral disambiguation through an open source and user-friendly graphical interface, published under DOI:10.6084/m9.figshare.26245895 [53], where a complete guide for installation and usage are also provided. Even though the dataset only contains XY8-$M$ simulations for NVs, it can be used for other sequences which have similar resonance spectra, such as the CPMG [10].
In the ambiguous resonances we studied here, the simulations model presented robust correlations with the experimental data. Numerically solving the system’s time dependent dynamics without the RWA has proven to be a more versatile and robust approach than semi-analytical models, recurring to perturbation theory. Although primarily focused with NVs, the developed model can be applied to any quantum system which satisfies the following three criteria - it can be described in terms of the density matrix formalism, it has a total Hamiltonian which can be divided in three parts as $H=H_{0}+H_{1}(t)+H_{2}(t)$ and it has a hermitian observable. Most color centers and many other quantum systems satisfy these conditions. Furthermore, the generality of $H_{1}(t)$ implies that the model can be extended to more elaborate DD sensing protocols as well.
@@ Acknowledgements.
Appendix A Experimental Setup and Data Acquisition
@@
The experiments were performed on a custom-made confocal microscope. The diamond samples are mounted on a custom made 3D stage movable with 10 $\mu$m precision. A diode laser with a wavelength of 518 nm (iBEAM-SMART-515-S, TOPTICA Photonics AG) was used for both continuous and pulsed optical excitation of an NV center. Emitted fluorescence was collected either with an air objective (MPlanApo N 50$\times$/0.95, Olympus) with NA = 0.95 or an oil objective (HCX PL APO 100$\times$/1.40, Leica; type F immersion liquid, Leica) with NA = 1.40 and detected by an avalanche photodiode (SPCM-AQRH-14, Excelitas). The objective is screwed on a piezo scanner (P-517, Physikinstrumente) with subnanometer resolution. The laser light was blocked from entering the detector by a 650 nm long pass filter (FELH0650, Thorlabs). Continuous MW field for ODMR spectra was generated by a microwave source (TGR6000, TTi) and the RF signal in Sec. IV.3 by a function generator (TGF4242 LXI, TTi). MW pulses for dynamical decoupling multipulse sequences were generated by an arbitrary waveform generator (AWG7122C, Tektronix), with maximum sampling frequency of 12 GSamples/s and memory of 64 MSamples. In both cases, generated MW fields were amplified (ZHL-16W-43+, MiniCircuit) and combined with the non-amplified RF using a bias tee (Mini Circuits, ZABT-2R15G+). Finally, they were transmitted to the NV centers by a copper (99.99%) wire with a diameter of 20 $\mu$m. The bias magnetic field was applied with a permanent magnet mounted on a holder with 4 degrees of freedom and with the maximum achievable magnetic field projection on NV center axis of 0.1 T. For time-resolved photon counting we used a time-correlated single photon counter (MCS6, FAST ComTec). Prior to all pulsed measurements, the $\pi$-pulse duration was measured with a Rabi sequence. The experimental value of the Rabi frequency was then used in the simulations. Lastly, laser pulses for state polarization and readout were applied for 3 $\mu$s and power from 0.1 to 0.5 mW, depending on the sample.
Appendix B Diamond Samples
@@
In Sec. IV.1, the effects of 15N coupling on Hahn echo were measured in a IIa diamond with a 12C overgrown layer of 90 nm, which was later implanted with 15N at energy 2 keV and annealed at 1000C in ultra-high vacuum. For the XY8-$M$ measurements, another shallow implanted sample was used with implantation energy 2 keV and dose $10^{9}$ 15N+/cm2, annealed at 10-7 mbar and 1000C. In Sec. IV.2, the 13C ambiguous resonance were studied in an electronic grade diamond grown by chemical vapor deposition, with natural abundance of 13C and 14N, in order to avoid 15N effects. Finally, the ambiguous resonances from finite pulses in Sec. IV.3 were measured in a thin electronic grade diamond plate with deeper 15N implantation (20-30 keV) and doses around $10^{9}$ 15N+/cm2, also annealed in ultra high vacuum at 1000C.
Appendix C Pearson Correlation Coefficient
@@
Figure 6: The XY8-2 of the ambiguous 15N resonance (Sec. IV.1) as a function of the pulse duration $T_{p}/t_{\pi}$ exemplifies the use of the Pearson coefficient, with r = (0.419, 0.858, 0.555, 0.685) for $T_{p}/t_{\pi}=$ (0.620, 0.826, 1.000, 1.215) respectively. At $T_{p}/t_{\pi}=1$ the simulated and experimental data have a poor correlation, but considering a small deviation, the experimental data can be better modeled by the simulation. The pulse error induces a high frequency modulation of the signal and variations in the peak intensities.
As discussed in Sec. III, experimental fluorescence intensity and simulated transition probability to the bright state ($m_{s}=0$) are related through a linear equation. Therefore, an appropriate metric to determine the correlation between both is the Pearson coefficient $r$ [40], which is the square root of the coefficient of determination $R^{2}$ for linear relations. It is defined as the ratio between the covariance of the two data sets and the product of their standard deviations, where one data set represents the experimental data and the other the simulation. Within this definition, $r$ is normalized from -1 to 1, with $r=0$ representing no linear dependency between experimental and simulated data, while $r=1$ indicates a perfect linear correlation.
To illustrate this, Fig. 6 shows simulations for XY8-2 of the 15N ambiguous resonances at $B_{0}=17$ mT for different pulse length deviations in regards to the actual $\pi$-pulse duration $T_{p}/t_{\pi}$ (Sec. IV.1). Compared to the actual experimental data, assuming $T_{p}/t_{\pi}=1$ results in a low correlation of $r=0.555$ and thus an inadequate simulation of the experimental data, whereas taking $T_{p}/t_{\pi}=0.826$ provides a better correlation of $r=0.858$. In this case, the pulse length deviation affects the relative intensity of each resonance peak keeping their $\tau$ fixed, while also inducing a high frequency modulation of the signal.
Appendix D 13C Hahn Echo Modulations
Figure 7: Hahn echo modulations from a 14NV coupled to a 13C nuclear spin, resulting in correlations between experimental and simulated data of $r=$ (0.792, 0.925) at $B_{0}=31$ mT and $r=$ (0.659, 0.806) at $B_{0}=18$ mT. Although both agree in regards to the modulation frequencies, the amplitude of the faster modulation is smaller in the experimental data, which could be due to some experimental error or unconsidered factor in the Hamiltonian.
@@
The ESEEM measured by a Hahn echo sequence at $B_{0}=31$ mT and $B_{0}=18$ mT for both $m_{s}$ transitions are shown in Fig. 7. As can be observed, the simulations and experimental data both present the same modulation frequencies, however the amplitude of the fast modulation is smaller in the experiments. This could be related to some experimental error or some unconsidered factor in the Hamiltonian model. Nonetheless, this discrepancy does not affect the study of the ambiguous resonances and as seen in Sec. IV.2, the calculated hyperfine matrix and the numerical model still reliably described them.
@@
Based on those experimental values of the modulation frequencies, different combinations of hyperfine coupling matrices between the NV and the 13C nuclear spin were tested in the total Hamiltonian $H_{0}+H_{2}$ from Eqs. 1 and 3. The eigenenergies of the Hamiltonians were calculated numerically, with the effective Larmor frequencies $\tilde{\omega}^{c}_{m_{s}}$ corresponding to the energy differences between the carbon nuclear spin states in each electron spin state $m_{s}$. Since no splittings could be observed in the two $m_{s}$ resonances, the ODMR spectrum sensitivity sets an upper bound for the $|A_{zz}^{c}|$ component of 4 MHz. Furthermore, the non-diagonal components of the 13C families are lower or of the same order as $|A_{zz}^{c}|$ [46], thus this bound was also assumed for the remaining components of $A^{c}$. Initially then, the six hyperfine elements were linearly varied in the interval from -4.0 MHz to 4.0 MHz with a large spacing of 0.5 MHz. Having found an approximate optimal region of the parameter space, each parameter was subsequently varied with spacing 0.01 MHz in this smaller range, totaling over $10^{7}$ possible combinations. Ultimately, the hyperfine matrix with the least square difference in units of MHz was found to be:
@@
Appendix E Pulse Errors
Figure 8: XY8-2 and RXY8-2 with external RF signal of $f_{0}=5$ MHz with MW pulse deviation of $\delta_{0}=20$ MHz or length error of $T_{p}/t_{\pi}=0.7$. The RF field is approximated by an effective hyperfine coupling with $A^{s}_{xz}=0.92$ MHz, resulting in $r=$(0.763, 0.907, 0.626) for XY8-2 and $r=$(0.639, 0.915, 0.423) in RXY8-2. The pulse errors lead to a degradation of the contrast and additional resonances, which are suppressed by the random phases.
Phase randomization can greatly increase robustness against pulse errors [49, 51]. In non-randomized XY8, small detunings in the MW frequency $\delta_{0}$ and deviations in the pulse length $T_{p}/t_{\pi}$ lead to a degradation of the resonance contrast and additional peaks over the $\tau$ range, which could represent another source of an ambiguous resonance for similar frequency RF fields or spins. To show this effect, XY8-2 and RXY8-2 with uncorrelated phases were measured and simulated with an applied RF field of $f_{0}=5$ MHz, taking $\delta_{0}=20$ MHz and $T_{p}/t_{\pi}=0.7$ separately (Fig. 8). Experimental parameters are $t_{\pi}=15.24$ ns and $B_{0}=34$ mT with $\theta_{0}=4.1^{\circ}$. Again, we observe a suppression of the spurious peaks in RXY8-2 and a further validation of the simulation model. In this case at low order $M=2$, the RF signal can be approximated by an effective hyperfine interaction with $A^{s}_{xz}=0.92$ MHz and $\gamma^{s}B_{0}=5$ MHz, as in Eq. 4. Whereas for the higher order XY8-12 from Sec. IV.3, the quantum mechanical approximation of the classical RF field already deviates from the experimental data. At the large detuning considered here $\delta_{0}=20$ MHz, the simulation shows smaller correlation with the experimental data. This evidences a limitation of the numerical model, while the reasons for it remain an open problem. Nonetheless, such large detuning might not be of practical significance for most experiments, apart from benchmarking the numerical simulations model.